# Using AI to Summarize US Presidential Campaign TV Advertisement Videos, 1952–2012

### v1

@@ Abstract
This paper introduces the largest and most comprehensive dataset of US presidential campaign television advertisements, available in digital format. The dataset also includes machine-searchable transcripts and high-quality summaries designed to facilitate a variety of academic research. To date, there has been great interest in collecting and analyzing US presidential campaign advertisements, but the need for manual procurement and annotation led many to rely on smaller subsets. We design a large-scale parallelized, AI-based analysis pipeline that automates the laborious process of preparing, transcribing, and summarizing videos. We then apply this methodology to the 9,707 presidential ads from the Julian P. Kanter Political Commercial Archive. We conduct extensive human evaluations to show that these transcripts and summaries match the quality of manually generated alternatives. We illustrate the value of this data by including an application that tracks the genesis and evolution of current focal issue areas over seven decades of presidential elections. Our analysis pipeline and codebase also show how to use LLM-based tools to obtain high-quality summaries for other video datasets.
@@
Despite this scholarly interest, two challenges remain that hinder further development of studies on US presidential campaign TV ads. First, the literature on ads has repeatedly emphasized the importance of analyzing how the content and style of ads have evolved over time [e.g., 13, 14]. However, such analysis is largely precluded by the fact that ad data are limited or unavailable for elections prior to 1996. Second, even when ads from recent elections are available, extracting relevant information from the videos for downstream analyses has required costly human hand-coding, limiting the scope of analyses to a narrow set of variables (see the Wesleyan Media Project [^1] http://mediaproject.wesleyan.edu/ and its predecessor Wisconsin Advertising Project [^2] http://elections.wisc.edu/wisconsin-advertising-project/). In part due to these limitations, scholars have focused primarily on analyzing small subsets of US presidential campaign TV ads [e.g., 15].
@@
Finally, we validate each step of our pipeline used to generate quality AI-generated summaries. Specifically, we validate our video preprocessing step, AI video transcription step, and AI video summarization step by comparing stratified samples of these outputs with human-generated outputs. Our main validation result is that human specialists find the quality of our LLM-generated ad transcripts and summaries to be at least equal to, and often superior to, the quality of transcripts and summaries composed by human research assistants across multiple key dimensions. Importantly, we also show that the high quality of our transcripts and summaries remains consistent across key sources of variation in the literature, including political party and time. This result and several others from the validation exercises outlined below will give applied researchers a great deal of confidence when they use both the raw ads and the associated summaries and transcripts for their own research.
Data Records
The Kanter presidential ad video collection
A former staffer for Adlai Stevenson, Julian P. Kanter started the Kanter political TV ad collection in $1956$ when he realized that political commercials were being discarded, or later taped over, after airing. He obtained the early film ads from radio and television studios, ad agencies, and the campaigns themselves. He later traded blank tapes in exchange for the already aired commercials. The collection moved to the University of Oklahoma in $1986$ when it outbid the Smithsonian to purchase approximately 25,000 radio and television ads from Julian Kanter for \$250,000 [18], including approximately 21,000 television ads ranging from 1950 to 1984 [19]. The archive has increased in size as Kanter and others continued collecting ads and political advertising firms donated their film and video. At the University of Oklahoma, the archive was originally housed in the Political Communications Center and is now part of the Carl Albert Congressional Research and Studies Center’s archival collections [20]. [^3] Users can search the full archive here, including numerous dataset collections: https://arc.ou.edu/.
@@
Although the collection has grown to more than $100{,}000$ commercials from campaigns up and down the ticket, we focus on the $9{,}707$ ads for the US presidency from $1952$-$2012$. Table 1 reports an overall summary of the ad collection. We do not make any claim that the collection contains the complete universe of ads (and it is not known which ads are missing), or that it is a representative sample of all ads. However, it is the most complete collection available in such a long time frame. Indeed, for a substantial number of items in the collection, the archive has the only known copy anywhere. To emphasize this point, we asked two undergraduate research assistants to check whether the ads in our dataset also appear in The Living Room Candidate [^4] https://www.livingroomcandidate.org/, which is the largest known collection of presidential ads (365 ads matching our $1952$-$2012$ time period) available online. Using a random sample of size $N=500$ from our collection, they found only $10.6$ percent of our ads were available on that website.
@@
In addition to the ads themselves, we also release a high-quality transcript for each ad. These transcripts serve three purposes. First, they permit practitioners to search the ad dataset for keywords of interest in order to identify ads that are relevant to their specific research topic. We provide online text search functionality on our project website https://campaignTVads.org/ (see Usage Notes below). Second, transcripts allow practitioners to study the dataset using novel methods from the recent renaissance in text-based research methods for social science [see e.g., 40, 41, 42, 43]. Finally, transcripts are also a key ingredient of our ad summarization workflow, as we describe below.
Summaries
We also provide a high-quality $50$-word summary of each ad designed to significantly accelerate the process of manually parsing the dataset. Historically, most research on political ads has focused on small subsets of ads, in no small part due to the high manual labor costs associated with viewing ads. Specifically, while each individual ad is relatively brief, our dataset contains more than $100$ hours of ads in total, which can prevent practitioners from conducting a thorough data exploration. Transcripts are helpful in this regard, but reading transcripts requires nearly as much time as watching videos and also ignores the rich visual information in the ads. Therefore, we also provide a high-quality $50$-word summary of each ad. These summaries offer an efficient synopsis of both audio and visual content, and each summary can be parsed by a human in approximately $4$ seconds.
@@ Methods
In this section, we discuss the proposed automated procedure that takes a historical campaign TV ad and produces a complete transcript and an efficient, high-quality $50$-word AI-generated summary designed to facilitate downstream research. We begin by describing how we preprocessed video files, as the original raw files contained extra segments at the beginning and end that were not part of the ad. We then show how to generate ad summaries using LLMs.
@@
We develop an automated trimming algorithm based on the following key observation; despite their diversity, nearly all political ads begin and end with a scene that contains video footage of people and/or audio of speech, such as voiceover or a fading in (or out) scene depicting a candidate or interviewer speaking. We leverage this observation by performing, on each video, (1) automated video transcription, (2) automated video person detection, and (3) automated video scene detection, using the Google Video Intelligence API [44]. [^5] We note that we used the Google Video Intelligence API for transcription during the preprocessing step only. Over the course of our project, we later determined that it performs as well as Whisper in terms of estimating the instant when speech begins (or ends), which is the goal of transcription as it is used in video preprocessing. However, the Google API is inferior to Whisper in terms of transcription word accuracy, which is the goal of Step 1 below. Researchers who wish to apply our methodology to their own datasets may skip using Google Video Intelligence API for transcription and simply use the Whisper beginning/ending timestamps instead, as they are computed in Step 1 below anyway. We then trim the beginning and end segments of the video file, yielding the video clip that spans from the beginning of the earliest scene that contains a person or spoken word to the end of the last scene that contains a person or spoken word.
@@
We accomplish this through two complementary strategies. First, we determine when important images are shown according to the ad’s narrator or speaker via to the transcription segments detected in Step 1 above. Specifically, we observe that the visual contents of a particular ad are almost always closely time-synced to its transcript. For example, the vast majority of scene changes correspond to transitions between speech segments, which we detected in Step 1 above. Moreover, most visually important imagery is onscreen around the middle of a narrator’s (or speaker’s) speech segment. Therefore, for each video, we extract a frame corresponding to the central moment of each speech segment. [^6] Our technique differs from [46] who randomly sample one frame for every 15 seconds of video, as well as one frame from the first and last two seconds of each ad. In the example shown in Table 2, we extract the frame that is depicted at the timestamp that falls in the middle of the phrase “I was born in …” (i.e. the frame at $(4.4-0)/2=2.2$ sec), another frame from the middle of the phrase “I remember that …” (i.e. the frame at $6.97$ sec), and so on.
@@
Technical Validation
@@
We compute two WER metrics for our validation. First, we compute the WER for each transcript and then take the mean over the transcripts. The transcripts have a mean WER of $0.0353$. Second, we also consider the overall WER, which we compute by concatenating the $120$ transcripts and calculating a single overall WER. This yields an overall WER of $0.0329$. In both cases, we conclude that our transcripts exhibit roughly $3$ errors per $100$ words. This means that our transcripts exhibit higher accuracy than standard benchmark transcriptions of virtually all standard transcription datasets except for ‘maximally clean’ high-definition speech data recorded in voiceover booths (e.g. transcriptions of LibriSpeech Clean data obtain WER that differs from ours by a margin of less than a single error in every $100$ words [see 45]).
@@
Here, $Y$ represents a Likert rating of a summary in one of the quality dimensions $d$, and $P(Y\leq k)$ denotes the cumulative probability of receiving a rating in category $k$ or lower. [^7] Note that the negative sign before the predictor effects means that larger values of $X\beta$ decrease the cumulative probability $P(Y_{d}\leq k)$ of a lower rating, shifting the probability mass toward higher ordinal categories. This matches our intuition that larger $X\beta$ corresponds to a better rating. The thresholds $\tau_{k}$ define the latent decision boundaries between ordinal categories. The terms in parentheses capture the main predictors of interest that estimate the treatment effect of AI summaries on ratings and control for covariates such as party and year. Finally, $u_{\text{Rater}}$ and $u_{\text{Ad}}$ terms are random intercepts that account for individual-level variation in rater tendencies and latent characteristics of the ad summaries, respectively. By including both random effects ($u$-terms) and covariates, the MEOL model allows us to properly account for the pairwise structure of the experiment, where each rater evaluates both the human- and AI-generated summary for the same ad, while also controlling for individual rater biases and latent ad characteristics that could influence ratings.
@@
Empirical Illustration: Tracking the Genesis and Evolution of Key Current Issues
We briefly illustrate the usage of our dataset by considering an application that measures the historical evolution of current key electoral issue areas. Understanding how key issue areas emerge is a central focus of the literature on political communication [13, 14, 46]. However, it has proven challenging to analyze this process in the absence of reliable data from elections prior to the 1996 election.
@@
Figure 4 plots the estimated proportions of topics for the $9$ issue-area-focused topics over the seven decades of ads in the dataset. While an in-depth analysis is beyond the scope of this paper, we find that the estimated proportions of foreign policy issues are inversely correlated with domestic issues related to the economy and taxation. For example, the postwar boom economy and tax issues dominate discourse in the $1950$’s, whereas these domestic policy topics cede to foreign policy issues that peak immediately following the Cuban Missile Crisis. The economy and tax issues experience resurgence during stagflation and again during Bill Clinton’s ‘it’s the economy, stupid’ campaign, ceding ground once more to foreign policy amidst the War on Terror, and finally experiencing an ‘issue area renaissance’ as America emerged from the $2008$ financial crisis. We also observe the nascence, ebb, and flow of the now-pervasive ‘crime’ issue area, beginning with the infamous ‘Horton weekend pass’ dog whistle that aired during Bush’s $1988$ campaign, and resurging in Bill Clinton’s infamous ‘superpredator’ and ‘three strikes’ marketing. We note that political campaign supply-side measures of content such as these offer an interesting point of comparison to standard opinion polls that only capture the ‘demand-side’ of political campaigns.
Figure 4: Evolution of key 2016-2024 election issues over time. We plot topic prevalence (measured via mean topic probability) over presidential elections from $1952$-$2012$. Topics focus on $9$ key issue areas, which represent the union of the top $3$ issues from each of the most recent $3$ presidential elections, $2016$-$2024$, plus two additional topics (taxation and crime) that reflect remaining issues among top-$3$ issue areas in terms of advertiser attention. Error ribbons represent $90$% credible intervals computed via the standard highest density interval method.
@@
Our project website (https://campaignTVads.org/) provides a convenient online interface designed to facilitate research by allowing practitioners to search and view videos that are relevant to their specific research questions.
@@
Downloading Bulk Videos, Transcripts, & Summaries
All presidential ads and associated transcripts and summaries are also available for bulk download on Harvard’s Dataverse (https://dataverse.harvard.edu/dataverse/ctap). We have organized the bulk data into subdirectories with one directory for each election, so that scholars interested in a particular subset of elections may bulk-download just that subset of interest.

### v2

@@ Abstract
This paper introduces the largest and most comprehensive dataset of US presidential campaign television advertisements, available in digital format. The dataset also includes machine-searchable transcripts and high-quality summaries designed to facilitate a variety of academic research. To date, there has been great interest in collecting and analyzing US presidential campaign advertisements, but the need for manual procurement and annotation has led many to rely on smaller subsets. We design a large-scale, parallelized, AI-based analysis pipeline that automates the laborious process of preparing, transcribing, storyboarding, and summarizing videos. We then apply this methodology to the 9,707 presidential ads from the Julian P. Kanter Political Commercial Archive. We conduct extensive human evaluations to show that these transcripts and summaries match the quality of manually generated alternatives. We illustrate the value of this data by including an application that tracks the genesis and evolution of current focal issue areas over seven decades of presidential elections. Our analysis pipeline and codebase also show how to use LLM-based tools to obtain high-quality summaries for other video datasets.
@@
Despite this scholarly interest, two challenges remain that hinder further development of studies on US presidential campaign TV ads. First, the literature on ads has repeatedly emphasized the importance of analyzing how the content and style of ads have evolved over time [e.g., 13, 14]. However, such analysis is largely precluded by the fact that ad data are limited or unavailable for elections prior to 1996. Second, even when ads from recent elections are available, extracting relevant information from the videos for downstream analyses has required costly human hand-coding, limiting the scope of analyses to a specific set of variables (see the Wesleyan Media Project [15] and its predecessor Wisconsin Advertising Project [16, 17]). In part due to these limitations, scholars have focused primarily on analyzing small subsets of US presidential campaign TV ads [e.g., 18].
Overview of the dataset: Ad videos, transcripts, and summaries, 1952-2012
@@
We validate each step of our pipeline used to generate quality AI-generated summaries. Specifically, we validate our video preprocessing step, AI video transcription step, and AI video summarization step by comparing stratified samples of these outputs with human-generated outputs. Our main validation result is that human specialists find the quality of our LLM-generated ad transcripts and summaries to be at least equal to, and often superior to, the quality of transcripts and summaries composed by human research assistants across multiple key dimensions. Importantly, we also show that the high quality of our transcripts and summaries remains consistent across key sources of variation in the literature, including political party and time. This result and several others from the validation exercises outlined below will give applied researchers a great deal of confidence when they use both the raw ads and the associated summaries and transcripts for their own research.
Finally, we also substantively validate and illustrate our dataset by considering an application that measures the historical evolution of current key electoral issue areas, as viewed through a keyword-assisted topic model estimated on the transcripts of all ads in the dataset. We show that the seven decades of data exhibit trends in campaign issues that correspond to historically recognized developments in campaign issue agendas.
History of the Kanter campaign advertisement video collection
A former staffer for Adlai Stevenson, Julian P. Kanter started the Kanter political TV ad collection in $1956$ when he realized that political commercials were being discarded, or later taped over, after airing. He obtained the early film ads from radio and television studios, ad agencies, and the campaigns themselves. He later traded blank tapes in exchange for the already aired commercials. The collection moved to the University of Oklahoma in $1986$ when it outbid the Smithsonian to purchase approximately 25,000 radio and television ads from Julian Kanter for \$250,000 [21], including approximately 21,000 television ads ranging from 1950 to 1984 [22]. The archive has increased in size as Kanter and others continued collecting ads and political advertising firms donated their film and video. At the University of Oklahoma, the archive was originally housed in the Political Communications Center and is now part of the Carl Albert Congressional Research and Studies Center’s archival collections [23]. Unfortunately, there are no presidential ads available in the dataset beyond the 2012 election. This is largely due to a resource constraint within the archive itself that limits active collecting and the reluctance of political consultants to donate commercials for current or potential future active candidates. We hope to add these recent ads to the dataset in the future and process them by applying the same techniques described in this paper. In the meantime, the missing ads limit what scholars are able to say about the last three elections in 2016, 2020, and 2024.
@@
The US presidential advertisement video dataset
Although the collection has grown to more than $100{,}000$ commercials from campaigns up and down the ticket, we focus on the $9{,}707$ ads for the US presidency from $1952$-$2012$. Table 1 reports an overall summary of the ad collection. We do not make any claim that the collection contains the complete universe of ads (and it is not known which ads are missing), or that it is a representative sample of all ads. However, it is the most complete collection available in such a long time frame. Indeed, for a substantial number of items in the collection, the archive has the only known copy anywhere. To emphasize this point, we asked two undergraduate research assistants to check whether the ads in our dataset also appear in The Living Room Candidate [24], which is the largest known collection of presidential ads (365 ads matching our $1952$-$2012$ time period) available online. Using a random sample of size $N=500$ from our collection, they found only $10.6$ percent of our ads were available on that website.
@@
In addition to the ads themselves, we also release a high-quality transcript for each ad. These transcripts serve three purposes. First, they permit practitioners to search the ad dataset for keywords of interest in order to identify ads that are relevant to their specific research topic. We provide online text search functionality on our project website https://campaignTVads.org/ (see Usage Notes below). Second, transcripts allow practitioners to study the dataset using novel methods from the recent renaissance in text-based research methods for social science [see e.g., 44, 45, 46, 47]. Finally, transcripts are also a key ingredient of our ad summarization workflow, as we describe below.
50-word summaries and visually representative ‘storyboard’ still frames
Historically, most research on political ads has focused on small subsets of ads, in no small part due to the high manual labor costs associated with viewing ads. Specifically, while each individual ad is relatively brief, our dataset contains more than $100$ hours of ads in total, which can prevent practitioners from conducting a thorough data exploration. Transcripts are helpful in this regard, but reading transcripts requires nearly as much time as watching videos and also ignores the rich visual information in the ads.
To address this problem, we take inspiration from a rich scholarly tradition of ‘campaign ad storyboards’—time-synced transcript segments associated with a visually representative selection of still frames extracted from each campaign ad (see, e.g., [48, 49]). Specifically, to supplement the ad transcripts, we develop a novel method to automatically extract a set of still frames that comprehensively represent the visual contents of each ad. We include these ‘storyboard’ still frames alongside each ad and transcript in our dataset. We then use the transcripts and these visually representative still frames to generate a high-quality $50$-word summary of each ad designed to significantly accelerate the process of manually parsing the dataset. These summaries offer an efficient synopsis of both audio and visual content, and each summary can be parsed by a human in approximately $4$ seconds.
Metadata
Finally, we also provide additional metadata about each ad. Our files include the candidate name, political party, election, ad title (if known), original video technology format (e.g., 16mm film), and ad duration.
@@ Methods
In this section, we discuss the proposed automated procedure that takes a historical campaign TV ad and produces a complete transcript, a visually representative selection of still frames, and an efficient, high-quality $50$-word AI-generated summary designed to facilitate downstream research. We begin by describing how we preprocessed video files, as the original raw files contained extra segments at the beginning and end that were not part of the ad. We then show how to generate ad summaries using LLMs.
Importantly, our summarization pipeline serves as readily-adaptable model. By providing the pipeline and data for all intermediate outputs, such as transcripts and key frames, we enable other researchers to generate their own custom-directed summaries or ad labels merely by adjusting the provided prompts and recomputing the final step of the pipeline, bypassing the significant cost and complexity of preprocessing, transcription, and frame extraction steps and their associated human validation workflows.
@@
We develop an automated trimming algorithm based on the following key observation; despite their diversity, nearly all political ads begin and end with a scene that contains video footage of people and/or audio of speech, such as voiceover or a fading in (or out) scene depicting a candidate or interviewer speaking. We leverage this observation by performing, on each video, (1) automated video detection of the timestamps of spoken words, (2) automated video person detection, and (3) automated video scene detection, using the Google Video Intelligence API [50]. As an aside, we note that researchers who wish to apply our methodology to their own datasets may skip using Google Video Intelligence API for spoken word timestamp detection and simply use the Whisper beginning/ending timestamps instead, as they are computed in Step 1 below anyway. We then trim the beginning and end segments of the video file, yielding the video clip that spans from the beginning of the earliest scene that contains a person or spoken word to the end of the last scene that contains a person or spoken word.
@@
Recent work has accomplished this by randomly sampling one frame for every 15 seconds of video, as well as one frame from the first and last two seconds of each ad [52]. However, our objective differs: rather than merely collecting a representative subset of frames, we aim to extract a frame sequence that comprehensively covers all visually relevant content in the ad.
We accomplish this through two complementary strategies. First, we determine when important images are shown according to the ad’s narrator or speaker via to the transcription segments detected in Step 1 above. Specifically, we observe that the visual contents of a particular ad are almost always closely time-synced to its transcript. For example, the vast majority of scene changes correspond to transitions between speech segments, which we detected in Step 1 above. Moreover, most visually important imagery is onscreen around the middle of a narrator’s (or speaker’s) speech segment. Therefore, for each video, we extract a frame corresponding to the central moment of each speech segment. In the example shown in Table 2, we extract the frame that is depicted at the timestamp that falls in the middle of the phrase “I was born in …” (i.e. the frame at $(4.4-0)/2=2.2$ sec), another frame from the middle of the phrase “I remember that …” (i.e. the frame at $6.97$ sec), and so on.
@@
Data Records
All presidential ads and associated transcripts, summaries, key frames, and metadata are available for bulk download on Harvard’s Dataverse at: https://dataverse.harvard.edu/dataverse/ctap  [56]. We have organized the bulk data into subdirectories with one directory for each election year so that scholars interested in a particular subset of elections may bulk-download just that subset of interest. The filenames for each transcript, summary, and ad key frame begin with the filename of the video to which they correspond. That video filename also matches a corresponding row in the dataset metadata file that includes the ad’s metadata. In the Usage Notes section below, we also describe our online interface designed to allow researchers to efficiently explore the data.
Technical Validation I: Data Accuracy, Consistency, and Balance versus Human Coders
@@
We compute two WER metrics for our validation. First, we compute the WER for each transcript and then take the mean over the transcripts. The transcripts have a mean WER of $0.0353$. Second, we also consider the overall WER, which we compute by concatenating the $120$ transcripts and calculating a single overall WER. This yields an overall WER of $0.0329$. In both cases, we conclude that our transcripts exhibit roughly $3$ errors per $100$ words. This means that our transcripts exhibit higher accuracy than standard benchmark transcriptions of virtually all standard transcription datasets except for ‘maximally clean’ high-definition speech data recorded in voiceover booths (e.g. transcriptions of LibriSpeech Clean data obtain WER that differs from ours by a margin of less than a single error in every $100$ words [see 51]).
@@
Here, $Y$ represents a Likert rating of a summary in one of the quality dimensions $d$, and $P(Y\leq k)$ denotes the cumulative probability of receiving a rating in category $k$ or lower. Note that the negative sign before the predictor effects means that larger values of $X\beta$ decrease the cumulative probability $P(Y_{d}\leq k)$ of a lower rating, shifting the probability mass toward higher ordinal categories. This matches our intuition that larger $X\beta$ corresponds to a better rating. The thresholds $\tau_{k}$ then define the latent decision boundaries between ordinal categories. The terms in parentheses capture the main predictors of interest that estimate the treatment effect of AI summaries on ratings and control for covariates such as party and year. Finally, $u_{\text{Rater}}$ and $u_{\text{Ad}}$ terms are random intercepts that account for individual-level variation in rater tendencies and latent characteristics of the ad summaries, respectively. By including both random effects ($u$-terms) and covariates, the MEOL model allows us to properly account for the pairwise structure of the experiment, where each rater evaluates both the human- and AI-generated summary for the same ad, while also controlling for individual rater biases and latent ad characteristics that could influence ratings.
@@
Technical Validation II: Tracing the Historical Emergence of Key Issues
We briefly validate and illustrate our dataset by considering an application that measures the historical evolution of current key electoral issue areas. A central focus of political communication research is understanding how such issues emerge and shift over time [13, 14, 63, 52]. This task not only demonstrates the dataset’s substantive relevance, but also serves as a validation check: do the trends it reveals correspond to historically recognized developments in campaign issue agendas? We find that our primary topics are generally consistent—though not perfectly aligned—with a well-known coding of subjects and themes that analyzes the transcripts from a smaller set of $895$ campaign advertisement from $1952$-$2000$ [63]. Such validation is especially valuable given the difficulty of analyzing electoral issue dynamics prior to the 1996 election due to limited data availability and high labor costs involved in coding more than a few hundred ads [63].
@@
Figure 4: Evolution of key 2016-2024 election issues over time. We plot topic prevalence (measured via mean topic probability) over presidential elections from $1952$-$2012$. Topics focus on $9$ key issue areas, which represent the union of the top $3$ issues from each of the most recent $3$ presidential elections, $2016$-$2024$, plus two additional topics (taxation and crime) that reflect remaining issues among top-$3$ issue areas in terms of advertiser attention. For presentation, we partition the issues into three plots corresponding to economic policy and social welfare, security and international affairs, and law and the courts. Error ribbons represent $90$% credible intervals computed via the standard highest density interval method.
Figure 4 plots the estimated proportions of topics for the $9$ issue-area-focused topics over the seven decades of ads in the dataset. While an in-depth analysis is beyond the scope of this paper, we find that the estimated proportions of foreign policy issues are inversely correlated with domestic issues related to the economy and taxation.
For example, consistent with the theoretical conjecture that incumbents should emphasize economic issues during periods of prosperity [63], ads in the 1950s often focus on the booming postwar economy and tax policy. These domestic topics gave way to foreign policy issues during periods when insurgent candidates attempted to draw attention to Cold War developments or when incumbents highlighted their opponents’ foreign policy positions. The 1964 campaign is illustrative: a spike in the foreign policy topic reflects Lyndon B. Johnson’s relentless critique—exemplified by the iconic “Daisy” ad—of Barry Goldwater’s escalatory support for deploying limited tactical nuclear weapons against communists in the immediate wake of the Cuban Missile Crisis and Johnson’s recent endorsement of the Nuclear Test Ban Treaty.
Economic and tax issues resurface during the stagflation era and again during Bill Clinton’s heresthetical [67] pivot to the “it’s the economy, stupid” strategy. These topics recede once more as foreign policy dominated the 2000s, particularly as George W. Bush focused his 2004 reelection campaign overwhelmingly on the War on Terror, favoring bellicose imagery and interventionist military rhetoric against Kerry’s more topically varied messaging that included a significant focus on healthcare reform (reflected in the spike in the healthcare topic in Fig. 4, top plot). In 2008, Barack Obama and John McCain maintained the central focus on terrorism and foreign affairs, though by then the discourse had largely shifted toward troop withdrawal and accountability. Obama’s campaign concurrently increased focus on economic topics as he refocused the contest through his “Wall Street versus Main Street” refrain and attempted to catch challenger McCain flat-footed on issues related to the contemporaneous onset of the financial crisis.
We also observe the nascence, ebb, and flow of the now-pervasive ‘crime’ issue area, beginning with the infamous ‘Horton weekend pass’ dog whistle that aired during Bush’s $1988$ campaign, and resurging in Bill Clinton’s infamous ‘superpredator’ and ‘three strikes’ marketing. We note that political campaign supply-side measures of content such as these offer an interesting point of comparison to standard opinion polls that only capture the ‘demand-side’ of political campaigns.
@@
In addition to bulk dataset access described in the Data Records section above, we also provide a project website (https://campaignTVads.org/) with a convenient online interface designed to facilitate research by allowing practitioners to search and view videos that are relevant to their specific research questions.